# Highfield, Southampton S017 1BJ, United Kingdom

### v1

@@ 5 The top-philic ALP as a portal to Dark Matter
In this section we consider the possibility that the top-philic ALP acts as a mediator between the SM and the dark-matter (DM) sector. Similar scenarios, for ALP mediators with other combinations of couplings with SM turned on, have previously been considered in, e.g., Refs. CidVidal:2018blh ; Acanfora:2023gzr ; Fitzpatrick:2023xks ; Dror:2023fyd (see also Refs. Arina:2016cqj ; Arcadi:2017wqi for the case of (pseudo)scalars). From the phenomenological point of view, it is an interesting possibility. On the one hand, since the ALP can feature an invisible decay, it opens up signatures with missing transverse momentum at the LHC. On the other hand, indirect limits obtained from the virtual exchange of a top-philic ALP, such as those from $t\bar{t}$ and $t\bar{t}t\bar{t}$ cross section measurements discussed in Sec. 4, remain applicable. Thus, both cases provide constraints that can be combined together.
@@ 5.1 The relic density constraint
As a final aside, note that in the regime $2m_{\text{DM}}<m_{a}$ we also expect stringent constraints from indirect detection, given the existing results from the Planck satellite data on thermal dark matter annihilating predominantly to $b\bar{b}$ Fermi-LAT:2015att ; Fermi-LAT:2016uux . Intriguingly, this is also the typical model and range of masses that could accommodate the galactic center excess (see e.g. Refs. Calore:2014nla ; PerezdelosHeros:2020qyt and references therein). A detailed investigation of direct and indirect detection constraints on this parameter space is left for future work.
@@ 6 Conclusions
In summary, our study shows that top-philic ALPs are quite elusive in the $[10,200]$ GeV mass range and that precision measurements at the LHC using top-quark final states can help significantly in improving the limits. We find that the visible and invisible scenarios can be constrained at the same level and that various channels can be used with similar sensitivities. An interesting direction that we leave for future work is to extend the mass window to fully cover the range up to the $2m_{t}$ threshold, where the ALP decays into (weak) gauge bosons become relevant and could could provide a powerful handle to constrain the model further.
@@ Acknowledgements
We thank Ilaria Brivio and Jesus Bonilla for many insightful discussions. DP and ST thank the Galileo Galilei Institute for Theoretical Physics for the hospitality during the completion of this work. FM is grateful to Alessandro Vicini for help with POWHEG. SB thanks Marco Gorghetto, Hyungjin Kim and Andrea Mitridate for useful discussions. SB and AM are supported in part by the Strategic Research Program High-Energy Physics of the Research Council of the Vrije Universiteit Brussel and by the iBOF “Unlocking the Dark Universe with Gravitational Wave Observations: from Quantum Optics to Quantum Gravity” of the Vlaamse Interuniversitaire Raad. SB, FM and AM are supported in part by the “Excellence of Science - EOS” - be.h project n.30820817. SB is in part supported by FWO-Vlaanderen through grant numbers 12B2323N. This work is supported by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy - EXC 2121 Quantum Universe - 390833306. KM is supported by an Ernest Rutherford Fellowship from the Science and Technologies Facilities Council (STFC), grant No. ST/X004155/1, and was previously supported by STFC grant No. ST/T000759/1.
@@ D.2 The $gg\rightarrow a\rightarrow t\bar{t}$ and $gg\rightarrow A\rightarrow t\bar{t}$ processes
We have cross-checked our implementation of the model ALPTopNLO_3flav UFO model presented in Appendix C, as also done for the $gq\rightarrow aq$ or $gq\rightarrow Aq$ process and the $qq\rightarrow ag$ or $qq\rightarrow Ag$ process, using the formulas collected in this Appendix. For the $gg\rightarrow a\rightarrow t\bar{t}$ and $gg\rightarrow A\rightarrow t\bar{t}$ process, the comparison is shown in Fig. 21 where the predictions from the formulae in Eqs. 79 and Eqs. 83 are plotted as lines, while the results obtained via MG5_aMC with the ALPTopNLO_3flav UFO model are shown as dots. In the left plot of Fig. 21 we show only the contribution of the interference of the one-loop diagram and the SM QCD one at tree level, Eqs. (82) and (85), while in the right plot we show only the contribution from the former diagram squared, Eqs. (81) and (84). In both plots the cross section is integrated of $\cos\theta$ and the azimuthal angle. As can be seen from the comparison of lines and dots, the results obtained via the ALPTopNLO_3flav UFO model are in agreement with the analytical formulae.
@@ Appendix F Limits from mono-jet searches on the invisible top-philic ALP
Figure 23: Bound on $f_{a}/c_{t}$ (GeV) for different $m_{a}$ (GeV). The strength of the bound depends on the kinematical region considered, modulated on the $x$-axis by the minimum $p_{\rm T}$ required for the jet.
@@
Out of these, we consider for each ALP mass the most stringent bound and report it in Tab. 7 and in Fig. 16. As expected from the behaviour of the $pp\rightarrow a+j$ cross-section illustrated in Fig. 4, the bound is relatively insensitive to the ALP mass. The low $p_{\rm T}^{\text{min}}$ region is not efficient in establishing a meaningful bound as we can see in Fig. 23. This can be understood by looking at how the cross-section of the $pp\rightarrow a+j$ process is distributed w.r.t. the transverse momentum of the jet in the top-philic ALP (see Fig. 6). The bound becomes more and more relevant when we increase the $p_{\rm T}^{\text{min}}$ value because the experimental searches become more constraining while the $p_{\rm T}$ differential distribution for $pp\rightarrow a+j$ decreases slowly due to the presence of the gluon contact interaction (note that the dip around $p_{\rm T}=1\,\textrm{TeV}$ is a fluctuation in the experimental data and is not related to the $pp\rightarrow a+j$ cross-section prediction).

### v2

@@ 5 The top-philic ALP as a portal to Dark Matter
In this section we consider the possibility that the top-philic ALP acts as a mediator between the SM and the dark-matter (DM) sector. Similar scenarios, for ALP mediators with other combinations of couplings with SM turned on, have previously been considered in, e.g., Refs. CidVidal:2018blh; Acanfora:2023gzr; Fitzpatrick:2023xks; Dror:2023fyd; Ghosh:2023tyz; Armando:2023zwz (see also Refs. Arina:2016cqj; Arcadi:2017wqi for the case of (pseudo)scalars). From the phenomenological point of view, it is an interesting possibility. On the one hand, since the ALP can feature an invisible decay, it opens up signatures with missing transverse momentum at the LHC. On the other hand, indirect limits obtained from the virtual exchange of a top-philic ALP, such as those from $t\bar{t}$ and $t\bar{t}t\bar{t}$ cross section measurements discussed in Sec. 4, remain applicable. Thus, both cases provide constraints that can be combined together.
@@ 5.1 The relic density constraint
As a final aside, note that in the regime $2m_{\text{DM}}<m_{a}$ we also expect stringent constraints from indirect detection, given the existing results from the Planck satellite data on thermal dark matter annihilating predominantly to $b\bar{b}$ Fermi-LAT:2015att; Fermi-LAT:2016uux. Intriguingly, this is also the typical model and range of masses that could accommodate the galactic center excess (see e.g. Refs. Calore:2014nla; PerezdelosHeros:2020qyt and references therein). Furthermore, also in the other mass regime, $m_{a}<m_{\rm DM}$, there might be indirect detection constrains from DM annihilation into SM pairs, even if these channels are subleading in setting the relic abundance (see e.g. Ref. Armando:2023zwz). A detailed investigation of direct and indirect detection constraints on this parameter space is left for future work.
@@ 6 Conclusions
In summary, our study shows that top-philic ALPs are quite elusive in the $[10,200]$ GeV mass range and that precision measurements at the LHC using top-quark final states can help significantly in improving the limits. We find that the visible and invisible scenarios can be constrained at the same level and that various channels can be used with similar sensitivities. An interesting direction that we leave for future work is to extend the mass window to fully cover the range up to the $2m_{t}$ threshold, where the ALP decays into (weak) gauge bosons become relevant and could provide a powerful handle to constrain the model further.
@@ Acknowledgements
We thank Ilaria Brivio and Jesus Bonilla for many insightful discussions. DP and ST thank the Galileo Galilei Institute for Theoretical Physics for the hospitality during the completion of this work. FM is grateful to Alessandro Vicini for help with POWHEG. SB thanks Marco Gorghetto, Hyungjin Kim and Andrea Mitridate for useful discussions. SB and AM are supported in part by the Strategic Research Program High-Energy Physics of the Research Council of the Vrije Universiteit Brussel and by the iBOF “Unlocking the Dark Universe with Gravitational Wave Observations: from Quantum Optics to Quantum Gravity” of the Vlaamse Interuniversitaire Raad. SB, FM and AM are supported in part by the “Excellence of Science - EOS” - be.h project n.30820817. SB is in part supported by FWO-Vlaanderen through grant numbers 12B2323N. This work is supported by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy - EXC 2121 Quantum Universe - 390833306. KM is supported by an Ernest Rutherford Fellowship from the Science and Technologies Facilities Council (STFC), grant No. ST/X004155/1, and was previously supported by STFC grant No. ST/T000759/1. ST is supported by a FRIA (Fonds pour la forma- tion ‘a la Recherche dans l’Industrie et dans l’Agriculture) Grant of the Belgian Fund for Research, F.R.S.-FNRS (Fonds de la Recherche Scientifique-FNRS)
@@ D.2 The $gg\rightarrow a\rightarrow t\bar{t}$ and $gg\rightarrow A\rightarrow t\bar{t}$ processes
We have cross-checked our implementation of the model ALPTopNLO_3flav UFO model presented in Appendix C, as also done for the $gq\rightarrow aq$ or $gq\rightarrow Aq$ process and the $qq\rightarrow ag$ or $qq\rightarrow Ag$ process, using the formulas collected in this Appendix. For the $gg\rightarrow a\rightarrow t\bar{t}$ and $gg\rightarrow A\rightarrow t\bar{t}$ process, the comparison is shown in Fig. 21 where the predictions from the formulae in Eqs. 80 and Eqs. 84 are plotted as lines, while the results obtained via MG5_aMC with the ALPTopNLO_3flav UFO model are shown as dots. In the left plot of Fig. 21 we show only the contribution of the interference of the one-loop diagram and the SM QCD one at tree level, [^19] In fact, in Sec. 4.5 we have considered only such contributions and not those from squared one-loop diagrams. Eqs. (83) and (86), while in the right plot we show only the contribution from the former diagram squared, Eqs. (82) and (85). In both plots the cross section is integrated of $\cos\theta$ and the azimuthal angle. As can be seen from the comparison of lines and dots, the results obtained via the ALPTopNLO_3flav UFO model are in agreement with the analytical formulae.
@@ Appendix F Limits from mono-jet searches on the invisible top-philic ALP
figureBound on $f_{a}/c_{t}$ (GeV) for different $m_{a}$ (GeV). The strength of the bound depends on the kinematical region considered, modulated on the $x$-axis by the minimum $p_{\rm T}$ required for the jet.
[…]
Out of these, we consider for each ALP mass the most stringent bound and report it in Tab. 7 and in Fig. 16. As expected from the behaviour of the $pp\rightarrow a+j$ cross-section illustrated in Fig. 4, the bound is relatively insensitive to the ALP mass. The low $p_{\rm T}^{\text{min}}$ region is not efficient in establishing a meaningful bound as we can see in Fig. F. This can be understood by looking at how the cross-section of the $pp\rightarrow a+j$ process is distributed w.r.t. the transverse momentum of the jet in the top-philic ALP (see Fig. 6). The bound becomes more and more relevant when we increase the $p_{\rm T}^{\text{min}}$ value because the experimental searches become more constraining while the $p_{\rm T}$ differential distribution for $pp\rightarrow a+j$ decreases slowly due to the presence of the gluon contact interaction (note that the dip around $p_{\rm T}=1\,\textrm{TeV}$ is a fluctuation in the experimental data and is not related to the $pp\rightarrow a+j$ cross-section prediction).